You are considering purchasing a $2,000, 20-year, 8% bond. Interest is paid semi-annually and comparable issues with similar risk are yielding 7%. If the bond has 12 years until maturity, what is the most you should pay for the bond today?
Answer #1
Price of this bond is the present value of cash flows of this bond.
Present value of coupon $       80.00 * 16.05837 = $              1,284.67
Present value of face value $ 2,000.00 * 0.437957 = $                  875.91
Present value of cash flows $              2,160.58
So, price of bond is $ 2,160.58 and I should pay $ 2,160.58 for this bond.
Working:
# 1 Semi annually coupon paid in cash = Face Value * Semi annual coupon rate
= 2000*8%*6/12
= $    80.00
# 2 Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.035)^-24)/0.035 i 7%/2 = 0.035
= 16.05837 n 12*2 = 24
# 3 Present value of 1 = (1+i)^-n
= (1+0.035)^-24
= 0.437957